OpenStudy (anonymous):

What is an irrational number besides a number that can't go on a number line, pi, and e?

OpenStudy (zzr0ck3r):

\[\sqrt{\text{prime number}}\]

OpenStudy (anonymous):

^ what he said...

OpenStudy (anonymous):

Are you sure?

OpenStudy (zzr0ck3r):

wait are you asking for another irrational number like pi and e?

OpenStudy (zzr0ck3r):

the wording of your question is odd

OpenStudy (anonymous):

OpenSessame, please leave... I will block you in fifty seconds.

OpenStudy (anonymous):

I just want to know what makes a number irrational.

OpenStudy (luigi0210):

Be nice @asdfasdf54514

OpenStudy (anonymous):

No, I'm rude.

OpenStudy (zzr0ck3r):

if it cant be written in the form\[\frac{integer}{integer}\]

OpenStudy (dan815):

when it cant be written as an integer

OpenStudy (zzr0ck3r):

so 1/2 is irrational dan?

OpenStudy (anonymous):

So, is .3333... Irrational?

OpenStudy (anonymous):

.333... is not irrational.

OpenStudy (dan815):

o true

OpenStudy (anonymous):

http://www.mathsisfun.com/irrational-numbers.html

OpenStudy (zzr0ck3r):

\[\frac{1}{3}=0.3333....\]

OpenStudy (anonymous):

THE GOLDEN RATIO!

OpenStudy (anonymous):

OpenSessame... Please... Stop... I only need ZZr0vk3r's help.

OpenStudy (anonymous):

Anyways, can it be written as irrational since the two integers of p and q are prime numberS?

OpenStudy (zzr0ck3r):

can what?

OpenStudy (anonymous):

1/3

OpenStudy (zzr0ck3r):

\[\frac{prime}{prime}\]is not an indication that something is irrational

OpenStudy (dan815):

irrational is when the numbers after the decimal point are just 'irrational' no pattern

OpenStudy (anonymous):

Oh yes! thank you!!! omg thank you so much!! it clicked now!! thank you guys!!

OpenStudy (anonymous):

TYSM NOW I CAN DO MY HOMEWORK WELL!!! THANK YOU TYTYTYTY!!

OpenStudy (zzr0ck3r):

I think you are trying find a tell all for irrational numbers, and i dont think there is. it simply means that it cane be put in a RATIO of integers

OpenStudy (zzr0ck3r):

cane=cant

OpenStudy (anonymous):

Yeah, I realized that! Thank you for helping me! I'm now a fan of yours for sure!

OpenStudy (zzr0ck3r):

\[\pi,e,\sqrt{3},\sqrt{2},....\]none of these numbers can be written in the form\[\frac{a}{b}\]where a, and b are integers

OpenStudy (zzr0ck3r):

np

OpenStudy (anonymous):

How come you're so smart?

OpenStudy (isaiah.feynman):

@zzr0ck3r 1 is not a prime number.

OpenStudy (zzr0ck3r):

im not, imagine someone asking you what 2+2 is, once you know it, you know it....

OpenStudy (zzr0ck3r):

i didnt say it was

OpenStudy (anonymous):

I'm studying mathematics theory, and my teacher is confusing me a lot.

OpenStudy (zzr0ck3r):

it was a counter argument...

OpenStudy (anonymous):

Do you have any tips for mathematics theory?

OpenStudy (isaiah.feynman):

@zzr0ck3r I see your point now.

OpenStudy (zzr0ck3r):

just practice, math theory really can mean anything. It depends on what you are doing..

OpenStudy (zzr0ck3r):

I like theory over applied

OpenStudy (anonymous):

Math Theory is going from pre algebra to differential calculus in a matter of 30 weeks.

OpenStudy (anonymous):

Thanks for your help!

OpenStudy (zzr0ck3r):

it all changes after you complete basic calc, difEQ, and linear algebra. things get very different once you get into calc analysis, real analysis, group theory/abstract algebra, differential geometry, ....these are all things you will do, and none of it will be anything like what you are doing now.

OpenStudy (anonymous):

Then why must I study the other maths if it is different?

OpenStudy (zzr0ck3r):

they are all the same thing, just looking through different glasses.

OpenStudy (zzr0ck3r):

not really but.....

OpenStudy (zzr0ck3r):

they are all realted....it would be like learning to throw a frisbie before you have learned to throw a ball. you dont need to learn how to throw the ball first, but it's probobally the best order to do it in.

OpenStudy (anonymous):

Oh, I understand, like how a bird learns to fly, but must first learn to dive out of a tree to impale its victim in an unforgivable manner.

OpenStudy (anonymous):

How come you're so wise?

OpenStudy (zzr0ck3r):

lol, im old:)

OpenStudy (isaiah.feynman):

@zzr0ck3r how old are you? and what level of math are you in???

OpenStudy (anonymous):

Old and Wise, like the man on top of the mountain. Hehehe :P

OpenStudy (zzr0ck3r):

im 34, im about to graduate with my bachelors in math

OpenStudy (anonymous):

i probabely smarter then zzrock asdfasd, he no smarter than me, i in college.

OpenStudy (anonymous):

LOL?

OpenStudy (zzr0ck3r):

next term im taking more real analysis, and more abstract algebra.

OpenStudy (anonymous):

I take math in college, it my majore, you take easy class zzrock

OpenStudy (anonymous):

What are you talking about?!

OpenStudy (zzr0ck3r):

lol you are trolling me on a math website?

OpenStudy (anonymous):

What troll? you no good at math man!

OpenStudy (anonymous):

I better than everone in evrthing!

OpenStudy (anonymous):

Zzr0ck3r, do you know him??

OpenStudy (anonymous):

I his brother I smarter in math ask me question insted

OpenStudy (zzr0ck3r):

hes just bored:)

OpenStudy (anonymous):

He left!

OpenStudy (anonymous):

I gotta go, take care man!

OpenStudy (dan815):

im going to be an astronaut

OpenStudy (dan815):

iam also going to be a doctor

OpenStudy (wolf1728):

asdfasdf54514 I can't believe the answer to this question has become so huge (yeah a lot of it is irrelevant information). An irrational number is one that cannot be put into the form of a fraction. 1/7 - although that is a repeating decimal .142857142857142857..., it is a rational number.

OpenStudy (dan815):

AND MAYBE a professional basketball player or baseball player maybe.. though

OpenStudy (isaiah.feynman):

@dan815 how old are you and what are you studying in school?

OpenStudy (dan815):

right now im in engineering physics

OpenStudy (dan815):

but im planning on taking more math in summer, complete a double major in maths and eng phys

OpenStudy (dan815):

starting 2nd year for physics in 5 more days T_T

OpenStudy (dan815):

OH MAN im not ready yet

OpenStudy (isaiah.feynman):

seems like you are still very young

OpenStudy (zzr0ck3r):

@wolf1728 you need to be careful with your wording its not that it cant be written in the form of a fraction as \[\frac{\pi}{2}\]is a fraction and is irrational the point is, that if the number is a real number and it cant be put in a ratio (fraction) where the numerator and denominator are integers then it is irrational.